### v2

@@ II-D Jaccard score
The size of the union is straightforward to define: it is the total number of points. In contrast, the intersection is a fuzzy notion due to points in the real and synthetic data generally being disjoint sets. Therefore, to define the intersection, we fit one Gaussian KDE model to the real data and a second Gaussian KDE to the synthetic data. We denote these real and synthetic KDEs $\hat{f}_{p}$ and $\hat{f}_{q}$, respectively. We take the intersection and union to be:
$p\cap q$ $=$ $\bigg{\{}p\in p\bigg{|}\frac{\hat{f}_{p}(p)}{\mathrm{max}\hat{f}_{s}}>0.1\bigg{\}}$ (2)
$p\cup q$ $=$ $\bigg{\{}q\in q\bigg{|}\frac{\hat{f}_{q}(q)}{\mathrm{max}\hat{f}_{r}}>0.1\bigg{\}}$ (3)
@@ IV-D Connections to entropy and diversity
IV-E Limitations and conclusions

### v3

@@ II-D Jaccard score
The size of the union is straightforward to define: it is the total number of points. In contrast, the intersection is a fuzzy notion due to points in the real and synthetic data generally being disjoint sets. Therefore, to define the intersection, we fit one Gaussian KDE model to the real data and a second Gaussian KDE to the synthetic data. We denote these real and synthetic KDEs $\hat{f}_{P}$ and $\hat{f}_{Q}$, respectively. We take the intersection to be:
$P\cap Q$ $=$ $\bigg{\{}p\in P\bigg{|}\frac{\hat{f}_{Q}(p)}{\mathrm{max}\hat{f}_{Q}}>0.1\bigg{\}}$ (2)
$\cup$ $\bigg{\{}q\in Q\bigg{|}\frac{\hat{f}_{P}(q)}{\mathrm{max}\hat{f}_{P}}>0.1\bigg{\}}$
@@ IV-D Connections to entropy and diversity
IV-E Use cases and exceptions
The general use case for the Eden score is to compare pairs of two-dimensional, i.e. pairwise, distributions. As we have shown, this makes it highly applicable for generative modeling, and for two-dimensional pairwise comparisons more generally. We propose that Eden could also be useful for comparing pairs of three-dimensional distributions, but caution that comparisons in still-higher dimensions could incur the curse of dimensionality.
The question arises of whether there two-dimensional cases where Eden might not be preferred. We believe there are at least two such cases. The first is where there is a strong prior on the two-dimensional data, for example when the expectation is that the data follows a simple line or curve. For example, suppose each of the two distributions is expected to lie on a simple straight line. In such cases, a regression-based score such as the correlation score can potentially still be useful; however, as Anscombe showed, it is still worth approaching with caution [13].
The second case where Eden might not be preferred is when the data configuration is very highly constrained. For example, in the extreme case where each feature is a binary variable (Boolean data), the space of possibilities consists of only four possible points: the coordinates (0,0), (0,1), (1,0), and (1,1). This constraint leaves no meaningful density regions as such. As a result, equidensity scores are unlikely to be of benefit. In such cases, an alternative approach is to treat the four coordinates as categoricals, and score agreement between the resulting two distributions accordingly. In contrast, Eden is for the general use case of complex and/or nonparameteric data defining meaningful densities in two dimensions. Potential additional use cases and limitations are left for future investigation.
IV-F Limitations and conclusions